# Benchmarking Multimodal LLMs on Recognition and Understanding over Chemical Tables

### v1

@@ 5.3 Analysis of Multimodal Input and Model Behavior
Case Study Analysis.
To evaluate the capabilities of MLLMs in handling descriptive and reasoning questions, we conducted a case study in Figure 6. The results reveal that MLLMs struggle significantly with complex reasoning tasks. Although proprietary models outperform open-source models in descriptive tasks, both exhibit notable limitations in accuracy and consistency, which highlights the challenges current MLLMs face in performing deep reasoning, underscoring the need for further advancements in multimodal reasoning capabilities.
@@ 6 Conclusion
In this work, we introduced ChemTable, a large-scale dataset and benchmark designed to evaluate multimodal large language models (MLLMs) on recognition and understanding tasks involving chemical tables. By curating over 1,300 real-world chemistry tables and annotating them with domain-specific metadata and question-answering tasks, ChemTable captured the multimodal, symbolic, and semantic challenges unique to the field. Our comprehensive evaluation revealed significant performance gaps between current MLLMs and human-level capabilities, particularly in domain-specific reasoning and molecular recognition. We believed that this dataset and benchmark would facilitate future advancements in multimodal scientific analysis and understanding.
@@ A.1 Effect of Question Hops on Answer Accuracy
To understand the impact of reasoning complexity on model performance, we conduct a fine-grained analysis of multi-hop question answering, using hop counts of 2, 3, and 4 as indicators of increasing logical depth. As shown in Table X, all models exhibit a monotonic decline in performance with increasing hop counts. For instance, Claude drops from 91.29% on 2-hop questions to 70.70% on 4-hop, while InternVL declines more sharply from 83.58% to 59.47%. This trend reflects a consistent rise in difficulty as models are required to perform more compositional and contextual reasoning over tabular data.
@@ A.2 Asymmetry in Condition-Yield Table Reasoning
Appendix B Correlation Between Response Length and Answer Correctness
@@ Appendix C Effect of Chain-of-Thought Reasoning
Appendix D Specifications of Annotation Format and Procedure for Table Recognition
@@ D.2 Fine-Grained Text Annotation Rules
  • Subscripts and Superscripts: Chemical text annotations follow the convention of using the caret symbol (^) for superscripts and the underscore (_) for subscripts.
@@ Appendix E Algorithm for Converting Annotations to HTML
6:    rowspan $\leftarrow$ 1 $+$ end_row -- start_row
7:    colspan $\leftarrow$ 1 $+$ end_col -- start_col
@@
10:             if row ==absent$=$ = start_row and col ==absent$=$ = start_col then
[…]
@@
10:             markup += <td rowspan $=rsp\text{ colspan}=csp\texttt{>}$
@@ F.2 LLM-Assisted Generation for Simple Reasoning
For numerical and statistically oriented reasoning tasks, we adopt a semi-automatic pipeline leveraging large language models (LLMs), specifically GPT-4.1. The question generation follows a prompt-based paradigm where we input the HTML representation of the table, the table image, and a specified reasoning type (e.g., comparison, summation) into a structured prompt template (see Section L for details). The model then outputs a set of QA pairs. The LLM is guided to focus on quantifiable patterns and basic statistical operations, such as:
@@ F.4 Domain-Specific Question Annotation: Function-Based QA
Agreement $\text{Agreement Rate}=\frac{|\mathcal{J}_{\text{human}}\cap\mathcal{J}_{\text{GPT}}|}{|\mathcal{J}_{\text{sampled}}|},$ (1)
where $\mathcal{J}_{\text{human}}$ and $\mathcal{J}_{\text{GPT}}$ denote the sets of instances labeled as correct by human annotators and GPT-4.1-nano, respectively, and $\mathcal{J}_{\text{sampled}}$ is the total set of sampled QA instances used for verification. The comparison between human annotations and GPT-4.1-nano’s binary classifications showed a high agreement rate of 96.8% overall. Agreement was particularly strong for descriptive and numerical tasks. These results confirm that GPT-4.1-nano provides a reliable and scalable approximation of human judgment for most evaluation scenarios in ChemTable, although human review remains valuable for borderline or high-complexity cases.
@@ Appendix H Screenshots of the Annotation Interface
Appendix I Implementation Details
@@
To ensure consistency and stability across different models during evaluation, we adopted a unified decoding configuration for all multimodal large language models. Specifically, the temperature was set to 0.0 and the nucleus sampling parameter (top-p) to 0.2, wherever supported. This configuration minimizes randomness and promotes deterministic outputs, which is essential for fair and reproducible evaluation across both table recognition and question answering tasks. This setup also facilitates meaningful comparisons between models under controlled conditions.
Appendix J Dataset Visualization by Image Type
@@
Appendix K Taxonomy of Question Types with Representative Cases
@@
How does the yield of 3p change with increasing reaction time? What is the trend in the third column? As k increases, what is the trend in the change of obs.$v_{a}-v_{b}$?
@@
Appendix L Prompt Templates
L.1 Table Recognition Prompt Templates
@@
L.2 Table Recognition Prompt Templates
@@
Appendix M Discussion of Limitations
This work focuses on evaluating general-purpose multimodal large language models on chemical tables but does not involve domain-specific pretraining or instruction tuning. As a result, current models may lack sufficient inductive bias for accurately interpreting chemical symbols, terminology, and molecular graphics. Future research could explore domain adaptation strategies such as finetuning (e.g., LoRA) to enhance models’ domain sensitivity.
Moreover, our evaluation isolates table images from surrounding textual context (e.g., experimental descriptions, chapter text), which often serve as essential semantic cues in scientific reading. Incorporating such context may improve performance on complex reasoning. Finally, the current benchmark emphasizes single-table understanding, whereas scientific synthesis often requires reasoning across multiple tables or documents. Extending the benchmark to multi-table or multi-document settings remains an important direction for future work.
Appendix N Potential Impacts
The ChemTable benchmark has the potential to positively influence both the scientific and machine learning communities by facilitating more accurate and structured extraction of domain-specific knowledge from chemical literature. This capability could accelerate AI-assisted scientific discovery, streamline the interpretation of experimental data, and improve accessibility for researchers engaging with complex multimodal content.
Nevertheless, there are possible negative societal impacts. Multimodal models applied to scientific documents may generate incorrect or misleading outputs, especially in high-stakes domains such as chemistry, where hallucinations or misinterpretations of molecular structures or reaction conditions could lead to unsafe conclusions. Moreover, if training data lacks diversity across chemical subfields or journal types, resulting models may exhibit biases or limited generalizability.
To mitigate these risks, we released the dataset under a permissive license and ensured that all data were sourced from publicly available and ethically compliant literature. We encourage future users to implement responsible usage policies, apply model interpretability tools, and evaluate models in safety-critical domains with domain-expert supervision.
Appendix O Copyright and Licensing

### v2

@@ 6 Conclusion
In this work, we introduced ChemTable, a large-scale dataset and benchmark designed to evaluate multimodal large language models (MLLMs) on recognition and understanding tasks involving chemical tables. By curating over 1,300 real-world chemistry tables and annotating them with domain-specific metadata and question-answering tasks, ChemTable captured the multimodal, symbolic, and semantic challenges unique to the field. Our comprehensive evaluation revealed significant performance gaps between current MLLMs and human-level capabilities, particularly in domain-specific reasoning and molecular recognition. We believe that this dataset and benchmark will facilitate future advancements in multimodal scientific analysis and understanding.
Ethics statement.
We confirm that this work aligns with accepted ethical standards in machine learning research. All datasets used in this study are derived from publicly available sources, and we carefully respect copyright and licensing requirements. Annotations were performed by qualified domain experts under fair working conditions. No personally identifiable or sensitive data were collected or used.
Reproducibility statement.
To support reproducibility, we provide detailed descriptions of the dataset construction process, annotation protocols, and experimental setups, including model configurations, hyperparameters, and evaluation metrics, in the main text and appendices. We also release the benchmark, scripts, and evaluation tools to facilitate replication and further research.
@@ A.1 Effect of Question Hops on Answer Accuracy
To understand the impact of reasoning complexity on model performance, we conduct a fine-grained analysis of multi-hop question answering, using hop counts of 2, 3, and 4 as indicators of increasing logical depth. As shown in Table 6, all models exhibit a monotonic decline in performance with increasing hop counts. For instance, Claude drops from 91.29% on 2-hop questions to 70.70% on 4-hop, while InternVL declines more sharply from 83.58% to 59.47%. This trend reflects a consistent rise in difficulty as models are required to perform more compositional and contextual reasoning over tabular data.
@@ A.2 Asymmetry in Condition-Yield Table Reasoning
Appendix B Correlation Between Response Length and Correctness
@@ Appendix C Effect of Chain-of-Thought Reasoning
Appendix D Specifications of Annotation Format and Procedure for TR
@@ D.2 Fine-Grained Text Annotation Rules
  • Subscripts and Superscripts: Chemical text annotations follow the convention of using the caret symbol (ˆ) for superscripts and the underscore (_) for subscripts.
@@ Appendix E Algorithm for Converting Annotations to HTML
6:  rowspan $\leftarrow$ 1 ++ end_row - start_row
7:  colspan $\leftarrow$ 1 ++ end_col - start_col
@@
10:     if row ==== start_row and col ==== start_col then
11:      table[row][col] $\leftarrow$ { "rowspan": rowspan,
@@
10:     markup += <td rowspan $rsp\text{ colspan}=csp\texttt{>}$
@@ F.2 LLM-Assisted Generation for Simple Reasoning
For numerical and statistically oriented reasoning tasks, we adopt a semi-automatic pipeline leveraging large language models (LLMs), specifically GPT-4.1. The question generation follows a prompt-based paradigm where we input the HTML representation of the table, the table image, and a specified reasoning type (e.g., comparison, summation) into a structured prompt template (see Section Q for details). The model then outputs a set of QA pairs. The LLM is guided to focus on quantifiable patterns and basic statistical operations, such as:
@@ F.4 Domain-Specific Question Annotation: Function-Based QA
$\text{Agreement Rate}=\frac{|\mathcal{J}_{\text{human}}\cap\mathcal{J}_{\text{GPT}}|}{|\mathcal{J}_{\text{sampled}}|},$ (1)
where $\mathcal{J}_{\text{human}}$ and $\mathcal{J}_{\text{GPT}}$ denote the sets of instances labeled as correct by human annotators and GPT-4.1-nano, respectively, and $\mathcal{J}_{\text{sampled}}$ is the total set of sampled QA instances used for verification. The comparison between human annotations and GPT-4.1-nano’s binary classifications showed a high agreement rate of 96.8% overall. Agreement was particularly strong for descriptive and numerical tasks. These results confirm that GPT-4.1-nano provides a reliable and scalable approximation of human judgment for most evaluation scenarios in ChemTable.
@@ Appendix H Screenshots of the Annotation Interface
Appendix I Dataset Distribution and Chemical Diversity
Figure 12: Year-wise distribution of papers in the dataset.
Figure 13: Temporal distribution of papers and extracted tables across the dataset.
The dataset was curated from leading chemistry journals between 2015 and 2024, ensuring both disciplinary relevance and temporal breadth. The majority of tables originate from Angewandte Chemie International Edition, Organic Letters, The Journal of Organic Chemistry, ACS Catalysis, and JACS, reflecting their central role in reporting experimental results. The year-wise distribution shows a steady increase in table usage, with recent years contributing the largest share, consistent with the growing trend of structured data reporting in chemical research. The detailed distribution can be found in Figures 13 and 13.
Table 7: Scaffold and reaction-type diversity statistics of the dataset.
Dimension Metric Value / Evidence
Scaffold diversity Unique Bemis–Murcko scaffolds 839
Scaffold-to-molecule ratio 0.208
Mean pairwise Tanimoto similarity 0.095
Top-3 scaffolds (share) Benzene 19.7% · No-ring 15.1% · Cyclohexane 10.9%
Reaction-type coverage Distinct reaction classes 15
Top-3 classes (share) C–C bond formation 16.4% · Oxidation 14.1% · C–Heteroatom bond 11.6%
Beyond bibliometric statistics, the dataset exhibits substantial chemical diversity. We identified 839 unique Bemis–Murcko scaffolds, yielding a scaffold-to-molecule ratio of 0.208 and a low mean pairwise Tanimoto similarity (0.095). The most frequent scaffolds include benzene (19.7%), acyclic frameworks (15.1%), and cyclohexane (10.9%), together covering less than half of the molecules. Reaction coverage spans 15 distinct classes, dominated by C–C bond formation (16.4%), oxidation (14.1%), and C–heteroatom bond formation (11.6%).
Taken together, these distributions highlight both the representativeness of the literature sources and the structural and functional diversity of the chemical space, making the dataset a robust testbed for benchmarking multimodal models in chemistry.
Appendix J Annotator Information and Consistency
The annotations for this study were performed by a team of chemistry experts with graduate-level education in the field. Their extensive training ensures a deep understanding of chemical terminology, experimental procedures, and domain-specific knowledge, which is crucial for the accurate interpretation and annotation of complex chemical tables.
To measure the consistency and reliability of the annotations, several quality control metrics were employed. The inter-annotator agreement (IoU) for cell boundaries reached 0.96, while the exact match accuracy for SMILES extraction was 0.99. Additionally, the inter-annotator agreement for cell content text was 0.94, further confirming the high consistency of the annotations across different annotators.
These quality control measures underscore the reliability of the annotated data set, making it suitable for subsequent analyzes and model evaluations in chemical table recognition and understanding tasks.
Appendix K Evaluating Domain-Specific and General Table Models on ChemTable Tasks
Figure 14: Performance comparison of domain-specific and general table models on ChemTable tasks.
This section evaluates the performance of domain-specific and general table models on chemistry-related tasks. Specifically, we compare ChemVLM (chemical domain-specific model) and Table-LLaVA 1.5 (general table model) in the context of chemical table understanding. The performance results across various question types are summarized in Figure 14.
The results indicate that general table models, such as Table-LLaVA, show limited transferability to chemistry tasks, primarily due to their lack of adaptation to the symbolic and multimodal nature of chemical tables. On the other hand, ChemVLM, which is specialized for the chemistry domain, performs better on certain tasks but still faces significant challenges in structured understanding and reasoning, especially with complex chemical data.
These findings highlight a substantial gap in model performance when it comes to understanding and processing chemical tables in scientific literature. While ChemVLM and Table-LLaVA perform well in their respective domains, they still fall significantly short of models like GPT-4.1 in addressing the unique challenges of chemical table comprehension. ChemTable serves as a benchmark that reveals these gaps, offering a realistic and challenging testbed for advancing model capabilities in chemical table recognition and reasoning.
Appendix L Implementation Details
@@ Appendix I Dataset Distribution and Chemical Diversity
To ensure consistency and stability across different models during evaluation, we adopted a unified decoding configuration for all multimodal large language models. Specifically, the temperature was set to 0.0 and the nucleus sampling parameter (top-p) to 0.2, wherever supported. This configuration minimizes randomness and promotes deterministic outputs, which is essential for fair and reproducible evaluation across both table recognition and question answering tasks.
Our human performance results are based on the annotations of five chemistry domain experts, all with graduate-level training. To ensure fairness, annotators were provided only with scratch paper and a basic calculator without advanced functions and were instructed to answer the given questions directly. We used a balanced assignment scheme in which each question was independently answered by three randomly selected annotators, and reported human performance as the average accuracy across their responses. Human performance is only reported for tasks that require chemical expertise or complex reasoning; purely descriptive element-level tasks are not annotated by humans.
Appendix M Qualitative Case Studies
We present several representative failure cases that reveal where current MLLMs struggle when processing chemical tables. These cases span four primary error types identified in our benchmark—(1) fine-grained recognition failures, (2) missing visual-style grounding, (3) domain-specific chemistry errors, and (4) multi-hop reasoning failures—and are further supplemented by an integrated example in the final, where multiple failure modes co-occur within a single real-world table. Together, these qualitative analyzes provide a more complete view of the limitations observed in both recognition and reasoning tasks.
M.1 Fine-Grained Recognition Errors
Figure 15: Case study of a fine-grained recognition failure on cell-level position retrieval: Gemini 2.5 Pro correctly OCRs and highlights the cell C(5) in a crystallographic coordinate table, but miscounts caption and header lines as data rows.
In Figure 15, we illustrate a typical fine-grained recognition error on the position-retrieval task. The table reports atomic coordinates from a crystallographic study, and the model is asked: *“What is the position of the cell with value ‘C(5)‘ in the picture?”* While Gemini 2.5 Pro correctly OCRs the entry ‘C(5)‘ and visually localizes it in the leftmost column, it still outputs an incorrect row index. The chain of thought shows that it sequentially numbers every line of the multi-line caption as table rows, then treats the header row as an additional row before starting to count the body entries. As a result, the true location “Row 10, Column 1” is misreported as “Row 14, Column 1,” even though the target cell itself is clearly identified. This case demonstrates that fine-grained structural alignment—mapping correctly recognized content to precise row–column indices—remains a key bottleneck, even when overall layout parsing and OCR quality appear to be strong.
M.2 Visual-Style Grounding Failures
Figure 16: Case study of visual-style grounding failure. The model forces an alignment between a query for a “yellow” column and a blue-shaded column, illustrating a failure in negative verification of visual attributes.
In Figure 16, we probe visual-style grounding with a descriptive QA prompt posed to GPT-5: “What is the header of the column with yellow color?” The ground-truth answer is “No cells matching this characteristic exist,” since the table uses only uniform grey shading to group yield columns and contains no yellow highlighting. GPT-5 nevertheless hallucinates a “colored” column, describing a shaded strip with 100% yields and confidently concluding that its header is 3a. This error shows that the model loosely maps any subtle background tint to the queried color instead of verifying the exact pixel-level style. Together with the large human–model gap on bold/color description in Table 4, this case indicates that even state-of-the-art MLLMs like GPT-5 still lack robust grounding of color and formatting cues in chemical tables: they can generate fluent style-aware explanations, but these explanations are often visually unsubstantiated.
M.3 Domain-Specific Symbol & Representation Challenges
Figure 17: Case study of a domain-specific notation failure: Claude-4.5-Sonnet misinterprets stereochemical footnote labels in an enantioselectivity table.
In Figure 17, we highlight how chemistry-specific notation can derail table understanding even for strong MLLMs like Claude-4.5-Sonnet. The table reports enantioselectivities for several cyclopropane diastereomers, with superscript footnotes tying each column to a particular stereoisomer. The model is asked: “What is the enantioselectivity (ee) for (1S,2R)-cyclopropane in the trans products when using the swMdo H64V/V68A catalyst?” The correct answer is that no such value is reported, because the relevant footnote actually refers to a different stereochemical assignment than the queried (1S,2R) trans product. Claude nonetheless treats footnote [e] as if it matched the target species, propagates this mislink through its chain-of-thought, and confidently returns “99% ee.” This case shows that current models still struggle to reliably bind domain-specific symbols and footnotes to their precise chemical meaning, and tend to hallucinate plausible numerical answers rather than recognizing when a requested quantity is formally undefined in the table, and refuse to answer this question.
M.4 Multi-Hop Reasoning Breakdowns
Figure 18: Case study of a multi-hop reasoning failure where Claude-4.5-Sonnet locates the correct row satisfying the yield constraints but outputs the base instead of the requested entry index.
In Figure 18, we stress-test multi-hop reasoning in Claude-4.5-Sonnet with a query that first requires locating the row where S4 and S5 have yields of 20% and 35%, and then returning the corresponding Entry index. The model successfully reads all numerical values, correctly identifies that only Entry 4 satisfies both constraints, and even states in its chain-of-thought that “Entry 4 matches both criteria exactly.” However, instead of outputting the requested index “4,” it reports the base used in that row, LDA (2.0 M in THF/EtPh/heptanes), effectively answering a different column than the one specified in the question. This mismatch between correct intermediate reasoning and the final prediction suggests that the bottleneck is not local retrieval or numerical comparison, but the final hop that maps a resolved row back to the correct schema field—an error pattern we observe repeatedly across multi-hop tasks in ChemTable.
M.5 Integrated Case Across Descriptive and Reasoning Tasks
Figure 19 provides a compact example illustrating how multiple failure modes can emerge within a single chemical table. For a simple descriptive query about the blue-labeled “Me” group, some models correctly interpret the colored annotation, while others misidentify it as the molecular formula of a nearby ligand, reflecting unstable grounding of domain-specific visual symbols. For a cell-level retrieval question, certain models also mislocate the target cell despite accurate OCR.
When the task shifts to chemically grounded reasoning—such as counting benzene rings in the ligand structures—all evaluated models produce large errors. This underscores the limitations highlighted: even strong MLLMs struggle with visual chemistry reasoning, especially in structures containing repeated or fused aromatic motifs. Overall, this example shows how symbolic misinterpretation, positional errors, and domain reasoning failures can co-occur within the same table context.
Appendix N QA Density and Category Distribution in ChemTable
We provide a detailed analysis of QA density in ChemTable to ensure that evaluation is not dominated by a small subset of tables or QA types. The final filtered split used for all experiments contains 9,886 QA pairs over 1,382 tables (7,344 descriptive + 2,542 reasoning QAs). The per-table QA density is moderate: the mean is 7.2 QAs per table (median 7, minimum 1, maximum 18, interquartile range 5–9). Only a small fraction of tables are very heavily annotated ($\approx 2\%$ of tables have $\geq 16$ questions), while over 80% of tables lie in the 5–15 QA range. The resulting Gini coefficient over the "QAs per table" distribution is 0.18, indicating a well-balanced and relatively uniform distribution. As shown in Table 8, the largest descriptive categories (Value Retrieval and Position Retrieval) each contribute only about 15–16% of all QAs. Reasoning-oriented categories such as Yield & Conditions, Multi-hop Retrieval, and Numerical Statistics are also well represented.
Table 8: Question–answer density and category statistics of the dataset.
Dimension Metric Value / Evidence
Overall QA density Tables with QA 1,382 tables
QA instances (total) 9,886 QAs (7,344 descriptive / 2,542 reasoning)
Mean / median QA per table 7.2 / 7
Min / max QA per table 1 / 18
25–75th percentile (QA per table) 5–9
Gini coefficient (QA density) 0.18 (lower is more uniform)
QA count per table 1–5 QAs per table 410 tables (29.7% of 1,382)
6–10 QAs per table 710 tables (51.4% of 1,382)
11–15 QAs per table 230 tables (16.6% of 1,382)
$\geq$16 QAs per table 32 tables (2.3% of 1,382)
Appendix O Dataset Visualization by Image Type
@@
Appendix P Taxonomy of Question Types with Representative Cases
@@
[…]
@@
Appendix Q Prompt Templates
Q.1 Table Recognition Prompt Templates
@@
Q.2 Table Recognition Prompt Templates
@@
Appendix R The Use of Large Language Models
We utilized large language models (LLMs) to assist and enhance the preparation of this manuscript. Specifically, LLMs were employed to improve clarity, grammar, and readability, while all conceptual, benchmark methodological, and experimental contributions are original and developed by the authors.
Appendix S Copyright and Licensing